OpenStudy (anonymous):

Nagaraj Nanjappa has a party-length sandwich 59in. long. He wants to cut it into three pieces so that the middle piece is 5 in. longer than the shortest piece and the shortest piece is 9in. shorter than the longest piece. How long should the three pieces be? (Help please, word problems are my weakest points :/)

OpenStudy (baldymcgee6):

Then make the word problem into a mathematical problem :)

OpenStudy (anonymous):

That's why I have trouble lol it.. word problems are my weakest points.. haha..

OpenStudy (anonymous):

SP=9 in MP=SP + 5 MP = 14 59 - 14 - 9 = 36 Thus the three answers are 9, 14, and 36 which equal 59.

OpenStudy (baldymcgee6):

so lets give this a try, x + (9+5) + 9 = 59

OpenStudy (anonymous):

now that i read 2 diff set ups im confused, xD..

OpenStudy (baldymcgee6):

x + (9+5) + 9 = 59 (unknown length) + (shortest piece + 5inches longer) +(shortest piece) = total sandwich (59inches)

OpenStudy (anonymous):

Okay, so then we'd have to subtract 9 from 50 and itself > x + (9+5) = 50?

OpenStudy (anonymous):

kinda lost.. sorry :x

OpenStudy (baldymcgee6):

Well in the equation, just solve for x: x + (9+5) + 9 = 59 x + (14) +9 = 59 x + 23 = 59 x = 59 - 23 x = 36 first part: 36 middle: 14 last part: 9 total: 59

OpenStudy (anonymous):

Oh I get it now! But can you give me a few tips to word problems so i can handle em with a bit more ease? What I have most trouble with is just forming the equation..

OpenStudy (baldymcgee6):

okay, for sure. Drawing pictures is always one method to solve a word problem efficiently, for example in this question.

OpenStudy (baldymcgee6):